OpenStudy (anonymous):

For f(x)= 1/x-5 and g(x)=x^2+2 find (f o g)(x) and (g o f)(6) This whole thing is still confusing to me... So I want to find f( g(x) ) correct? for the (f o g)(X), but I'm confused as to how I do it with x versus a regular number? Btw, If this is super obvious, sorry that I'm asking, I just want to understand it all.

OpenStudy (the_fizicx99):

f o g(x)= 1/(x^2 + 2) -5 = 1/x^2 -3 g o f(6)= (1/6-5)^2 + 2

OpenStudy (the_fizicx99):

I believe this is how its done, its been a while since I've done this lol

OpenStudy (opcode):

Composite functions, eek.

OpenStudy (the_fizicx99):

Wot? wrong?

OpenStudy (opcode):

I don't know I didn't check, seems right though. I dislike composite functions... (What's the point, to them...)

OpenStudy (the_fizicx99):

>.> yesh prolly true